A ship with a faulty engine sailed 7hrs over a period.of.2days covering 59km.Average speed on the first day was 5km per hour less than the 2nd day but sailed 3hrs longer than on the second day.What is the average speed of 2nd day?
Answer by ankor@dixie-net.com(22740)

A ship with a faulty engine sailed 7hrs over a period.of.2days covering 59km.
Average speed on the first day was 5km per hour less than the 2nd day
but sailed 3hrs longer than on the second day.
What is the average speed of 2nd day?
:
Let d = no. of hrs the 2nd day, then
(d+3) + d = 7 hrs
2d = 7 - 3
2d = 4
d = 2 hrs on the 2nd day
then
2 + 3 = 5 hrs on the 1st day
:
let s = av speed on the 2nd day
then
(s-5) = ave speed on the 1st day
:
1st day dist + 2nd day dist = 59
5(s-5) + 2s = 59
5s - 25 + 2s = 59
7s = 59 + 25
7s = 84
s = 84/7
s = 12 mph on the 2nd day
:
:
see if that checks out
1st day dist: 5(12-5) = 35 mi
2nd day dist: 2 * 12 = 24 mi
--------------------------------
two day total distance: 59 mi